### v2

# LMM-PCQA: Assisting Point Cloud Quality Assessment with LMM

Abstract.
Although large multi-modality models (LMMs) have seen extensive exploration and application in various quality assessment studies, their integration into Point Cloud Quality Assessment (PCQA) remains unexplored. Given LMMs’ exceptional performance and robustness in low-level vision and quality assessment tasks, this study aims to investigate the feasibility of imparting PCQA knowledge to LMMs through text supervision. To achieve this, we transform quality labels into textual descriptions during the fine-tuning phase, enabling LMMs to derive quality rating logits from 2D projections of point clouds. To compensate for the loss of perception in the 3D domain, structural features are extracted as well. These quality logits and structural features are then combined and regressed into quality scores. Our experimental results affirm the effectiveness of our approach, showcasing a novel integration of LMMs into PCQA that enhances model understanding and assessment accuracy. We hope our contributions can inspire subsequent investigations into the fusion of LMMs with PCQA, fostering advancements in 3D visual quality analysis and beyond. The code is available at https://github.com/zzc-1998/LMM-PCQA.
Large multi-modality model, Point cloud quality assessment
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3680946isbn: 979-8-4007-0686-8/24/10ccs: Human-centered computing Visualization design and evaluation methodsccs: Computing methodologies Artificial intelligence
Figure 1. Inspired by the impressive quality evaluation ability of LMM on 2D media, we are the first to explore the quality representation potential of LMM on 3D point clouds.
1. Introduction
Point clouds are increasingly used across diverse real-world scenarios, including virtual/augmented reality (Park et al., 2008; Liu et al., 2020; Gu et al., 2019), autonomous vehicles (Cui et al., 2021), and video post-production (Mekuria et al., 2016a). This surge is attributed to their adeptness in three-dimensional representation. Consequently, significant research efforts have been channeled towards enhancing high-level areas like point cloud classification (Grilli et al., 2017; Ku et al., 2018; Wang and Jia, 2019; Vora et al., 2020; Xie et al., 2020; Yoo et al., 2020; Chen et al., 2020), detection (Cui et al., 2021), and segmentation (Cheng et al., 2021; Liu et al., 2022a). Meanwhile, as a key component for ensuring point cloud quality, point cloud quality assessment (PCQA) has seen comparable advancements as well during the last decade. PCQA’s objective is to appraise point clouds’ visual quality, a pivotal factor for refining simplification and compression strategies in practical applications (Fan et al., 2022a), and to elevate the Quality of Experience (QoE) for end-users. Generally, PCQA methods are divided into three categories, depending on their reliance on reference point clouds: Full-Reference PCQA (FR-PCQA), Reduced-Reference PCQA (RR-PCQA), and No-Reference PCQA (NR-PCQA) respective. The availability of pristine reference point clouds is often limited in real-world applications, emphasizing the demand for NR-PCQA approaches, which is why our research primarily focuses on NR-PCQA.
There are already some cutting-edge studies that have begun applying Large Multi-Modality Models (LMMs) to low-level vision and quality assessment fields (Wu et al., 2024a, b, 2023; You et al., 2023; Huang et al., 2024), achieving notable success. LMMs demonstrate highly competitive performance and robustness in these tasks. However, the main focus of these studies has remained on two-dimensional (2D) media such as images and videos, while no research has explored the possibility of applying LMMs to three-dimensional (3D) media like point clouds. It is known that both 2D and 3D media exhibit similar distortions, i.e., blur and noise. Given the robust quality perception of LMMs in 2D, we can hold the hypothesis that LMM also has significant quality perception abilities in 3D point clouds. Hence, investigating the application of LMMs for point cloud quality assessment is not only valuable but also meaningful, reflecting their established visual perception strengths. Therefore, in this study, we carry out a novel method named LMM-PCQA to provide an interesting solution for handling the PCQA problem with the assistance of LMM.
First, we treat the point clouds as sequences of projections to enable LMM to perceive the point cloud visual quality. Afterward, we try to teach LMM about the quality alignment between the predefined 5-level qualitative adjectives (i.e., excellent, good, fair, poor, bad) and point cloud projections. Specifically, we employ the existing PCQA databases to provide the necessary knowledge, where the quality labels are transformed into corresponding qualitative adjectives. Then we specially design a prompt structure to produce the question-answer pairs, which are composed of the question ‘Rate the quality of the point cloud from the projections [img1],[img2],[img3],[img4],[img5],[img6]’ and the answer ‘The quality of the point cloud is excellent/good/fair/poor/bad’. These question-answer pairs can be utilized to teach LMM PCQA knowledge during the fine-tuning stage. After the instruction tuning, we can expect the trained LMM to give the predicted [SCORE_TOKEN] with the same prompt structure (the qualitative adjectives position is left blank for LMM response), which is shown in Fig. 2. The predicted [SCORE_TOKEN] can be recognized as a probability map to the qualitative adjectives, and we convert it into 5-level probabilities as the LMM evaluation results.
Secondly, to address the potential insensitivity to geometric distortions (e.g., compression, downsampling) when only analyzing projections, we propose the extraction of multi-scale structural features. This approach enhances the LMM-PCQA’s holistic comprehension of point cloud visual quality. The point clouds are converted into quality-aware structural domains, a technique validated for effective quality feature extraction in prior research (Alexiou et al., 2021; Zhou et al., 2023; Zhang et al., 2022b). We modify the scale parameters in the k-nearest neighbors (k-NN) algorithm to offer a multi-scale perspective, aligning with the human vision system’s perception mechanism. Subsequently, key statistical parameters are used to quantify structural distortions within these domains. Finally, we combine the LMM’s evaluative results and the structural features, utilizing support vector regression (SVR) to derive the quality values. The experimental outcomes affirm that our LMM-PCQA model is on par with, or surpasses, current leading PCQA methods.
By carrying out LMM-PCQA, the contributions of this paper can be summarized as follows:
  • We are the first to employ LMM for PCQA tasks. We design a novel prompt structure to enable the LMM to perceive the point cloud visual quality. The existing PCQA databases are converted into question-answer pairs, which are then used to inject PCQA knowledge into LMM.
  • We propose the extraction of multi-scale structural features. By processing point clouds into multi-scale domains, we quantify the geometry distortions via key statistic parameters estimation, which helps LMM gain a more comprehensive understanding of point cloud visual quality.
  • LMM-PCQA demonstrates exceptional performance across various PCQA databases. The ablation study and cross-database evaluations further validate the logical design of LMM-PCQA and its robust generalization capabilities.
2. Related Works
2.1. LMM for Quality Assessment
Recent studies are exploring LMMs in visual quality assessment. X-IQE (Chen, 2023) uses them to evaluate text-to-image generation via a chain of thoughts strategy. Q-Bench (Wu et al., 2024a) introduces a binary softmax method for LMMs to produce quantifiable scores by assessing tokens good and poor. Q-Instruct (Wu et al., 2024b) improves this by fine-tuning LMMs through text-based questioning for specific visual queries. Q-Align (Wu et al., 2023) adopts human-like evaluation mechanisms, enhancing visual quality scoring.
Figure 2. The framework of the proposed method.
2.2. PCQA Development
In the early stages of PCQA, the MPEG group developed Full-Reference PCQA (FR-PCQA) methods like p2point (Mekuria et al., 2016b) and p2plane (Tian et al., 2017) for point cloud quality assessment, and later introduced a color-based PSNR-yuv method (Torlig et al., 2018). Despite challenges in addressing complex distortions, advanced FR-PCQA metrics like PCQM (Meynet et al., 2020), GraphSIM (Yang et al., 2020b), and PointSSIM (Alexiou and Ebrahimi, 2020) were developed to integrate structural features for enhanced performance.
Building on progress in no-reference image and video quality assessment (NR-I/VQA) (Zhang et al., 2015a; Liu et al., 2015; Gu et al., 2017; Zhang and Liu, 2022; Zhang et al., 2022a, 2021a), various NR-PCQA methods have evolved. Techniques include using CNNs for quality regression with handcrafted features (Chetouani et al., 2021), multi-view projections for feature extraction (Liu et al., 2021b; Zhang et al., 2022b), and direct prediction with sparse CNNs (Liu et al., 2022c). Innovative approaches also involve converting point clouds to videos for VQA (Fan et al., 2022b; Zhang et al., 2022c), employing structure-guided resampling (Zhou et al., 2022), integrating multi-projections for simpler feature extraction (Zhang et al., 2023b), adapting image quality assessment methods for point cloud rendering (Yang et al., 2022), and using non-local geometry and color gradient models for quality estimation (Wang et al., 2023). Recent frameworks like MM-PCQA (Zhang et al., 2023a) and pmBQA (Xie et al., 2023) leverage multi-modal learning for improved PCQA. Emerging research explores text modality’s potential in quality assessment learning.
3. Proposed Method
The framework of the proposed method is briefly illustrated in Fig. 2, which includes the LMM evaluation module, the structural feature extraction module, and the quality regression module.
3.1. LMM Evaluation
3.1.1. Projection Acquisition
Consider a colored point cloud $\mathbf{P}={(p_{i},c_{i})}_{i=1}^{N}$, where $p_{i}\in\mathbb{R}^{1\times 3}$ represents the single point consisting of geometry coordinates, $c_{i}\in\mathbb{R}^{1\times 3}$ represents the RGB color attributes, and $N$ denotes the total count of points. We then adopt the conventional cube-like viewpoints configuration, which is inspired by the projections approach discussed in (Torlig et al., 2018; Alexiou et al., 2019). As illustrated in Fig. 2, six orthogonal viewpoints are utilized, each mapping to one of the cube’s six faces for generating the projections. For a point cloud $\mathbf{P}$, the rendering process is derived as:
(1) $\mathbf{I}=\psi(\mathbf{P}),$
[…]
where $\mathbf{I}$ represents the set of the 6 rendered projections and $\psi(\cdot)$ denotes the Open3D-based (Zhou et al., 2018) projections capturing process.
3.1.2. How to inject PCQA knowledge into LMMs ?
LMM has shown competitive performance in 2D image/video quality assessment (Wu et al., 2023, 2024a, 2024b), therefore it is feasible to apply LMM for PCQA tasks by taking the point cloud as a sequence of projections. Then it is vital to solve the core problem of PCQA Knowledge Injection. Following the common approaches of training LMMs (Ye et al., 2023a; Liu et al., 2023), it is natural to come up with the solution of instructing the LMM with question-answer pairs regarding PCQA issues. Thus we carry out the specific prompt structure as follows:
-How would you rate the quality of the point cloud from the projections?¡—img1—¿¡—img2—¿…
-The quality of this point cloud is [QA(mos)].
where ¡—img1—¿¡—img2—¿… stands for the image set of projections and QA(s) is the qualitative adjective of the point cloud which can be obtained from the mean opinion score (MOS) corresponding to the point cloud. Afterward, we can use the designed question-answer pairs to teach LMM PCQA knowledge.
3.1.3. Transformation from MOS to quality rating.
In daily experiences, humans often give feedback using qualitative adjectives (such as good, bad, superb) instead of numerical ratings (like 9.2, 2.5, 7.1). Hence, implementing visual scoring activities with level-based ratings taps into this natural tendency of humans (to offer qualitative adjectives). Similarly, the perception and expression of LMM are akin to humans, which have a better understanding and perception of qualitative adjectives. Therefore, converting MOS into corresponding qualitative adjectives for its learning is more intuitive than having it learn directly from numbers. Specifically, the transformation of MOS can be achieved by evenly splitting the range from the highest score ($\mathrm{M}$) to the lowest score ($\mathrm{m}$) into five unique intervals, with scores in each interval designated as corresponding quality levels:
(2) ${QA(mos)}\!=\!w_{i}\text{ if }\text{m}+\frac{i-1}{5}\times\mathrm{(M-m)}\!<\!mos\!\leq\!\mathrm{m}+\frac{i}{5}\times\mathrm{(M-m)},$
[…]
Figure 3. Illustration of the LLM evaluation pipeline. The point clouds with MOSs are transformed into question-answer pairs for LMM tuning. The LMM evaluation results can be obtained as the set of the probabilities to the predefined qualitative adjectives.
3.1.4. Obtaining evaluation results via LMM inference.
After training, we can get the evaluation results with the same prompt structure and get the response [SCORE_TOKEN] via LMM inference. The [SCORE_TOKEN] can be recognized as a log probability map to the qualitative adjectives. Then we can compute the final probabilities to the 5-level qualitative adjectives from the corresponding log probabilities via softmax as the LMM evaluation results:
(3) $F_{L}=\{\frac{e^{\alpha_{i}}}{\sum_{j=1}^{5}e^{\alpha_{j}}}\}_{i=1}^{5},$
where $\alpha_{i}$ indicates the log probability of $i$-th qualitative adjectives and $F_{L}$ represents the LMM evaluation results which consist of 5 probabilities after softmax.
3.2. Structual Feature Extraction
The inadequacy of single-modality information from projections for point cloud quality evaluation has been established (Zhang et al., 2023a; Xie et al., 2023). Therefore, our approach enhances the accuracy of LMM evaluation results (projections only) by integrating geometric structural features, aiming for a more detailed and accurate assessment.
3.2.1. Structural Domain
Given the point cloud $\mathbf{P}={(p_{i},c_{i})}_{i=1}^{N}$, the neighborhood $P_{Nbi}$ of each point $p_{i}$ can be obtained utilizing the k-nearest neighbors (k-NN) algorithm:
(4) $\quad\quad\mathbf{P_{Nb}}=\mathop{\boldsymbol{\rm{KNN}}}\limits(\mathbf{P}),$
$Dist(p,q)$ $=\!\sqrt{\left(p_{x}\!-\!q_{x}\right)^{2}\!+\!\left(p_{y}\!-\!q_{y}\right)^{2}\!+\!\left(p_{z}\!-\!q_{z}\right)^{2}},$
where $N$ represents the total count of points within the point cloud, the term $\mathbf{P_{Nb}}$ denotes the collection of neighborhoods, while $\boldsymbol{\rm{KNN}}(\cdot)$ signifies the function of the k-nearest neighbors algorithm. The distance between points $p$ and $q$ is calculated using the Euclidean distance, expressed as $Dist(p,q)$. Given the neighborhood set $P_{Nbi}$ of point $p_{i}$, we can define the covariance matrix $C_{i}$ for each point $p_{i}$, which is characterized by its 3D geometric coordinates:
(5) $C_{i}=$ $\frac{1}{K}\sum_{j=1}^{K}({p}_{n_{j}}-\hat{p})({p}_{n_{j}}-\hat{p})^{\top},$
$\{{p}_{n_{1}},\cdots,{p_{n_{K}}}\}\in P_{Nbi},$
where the term $K$ denotes size of neighborhood set $P_{Nbi}$, ${p}_{n_{j}}$ is the $j$-th neighboring point in $P_{Nbi}$, $\hat{p}$ is the centroid of this neighborhood, ${p}_{n_{j}}$ and $\hat{p}$ are vectors with dimensions $\mathbb{R}^{3\times 1}$, while $C_{i}$ is a matrix with dimensions $\mathbb{R}^{3\times 3}$. Consequently, the eigenvectors for the covariance matrix $C_{i}$ can be derived as follows:
(6) $C_{i}\cdot{v}_{l}=\lambda_{l}\cdot{v}_{l},l\in\{1,2,3\},$
where ($\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$) stand for the eigenvalues and (${v}{1}$, ${v}{2}$, ${v}{3}$) represent the respective eigenvectors, with $\lambda{1}>\lambda_{2}>\lambda_{3}$. Consequently, we derive three eigenvalues for each point $p_{i}$ within the point cloud $\mathbf{P}$. Then we can compute the linearity and planarity structural domains of the point cloud as:
(7) ${Lin(p_{i})}$ $=\frac{\lambda_{1}-\lambda_{2}}{\lambda_{1}},$
${Pla(p_{i})}$ $=\frac{\lambda_{2}-\lambda_{3}}{\lambda_{1}},$
where $Lin(p_{i})$ and $Pla(p_{i})$ represent the linearity and planarity values for point $p_{i}$. The chosen structural domains (linearity, planarity) have been demonstrated to exhibit a strong correlation with geometric visual losses, such as compression and downsampling, and have been extensively utilized in numerous PCQA tasks (Alexiou et al., 2021; Zhou et al., 2023; Zhang et al., 2022b).
3.2.2. Multi-scale Perception
The multi-scale nature of point cloud visual perception has been noted in the literature (Zhang et al., 2021b). To account for this, we compute structural domains across various scales by varying the scale parameter $k$ in the $\boldsymbol{\rm{KNN}}(\cdot)$ process. This approach allows us to derive the multi-scale structural domains as follows:
(8) $\mathcal{D}_{k=k_{ms}}=\mathcal{S}(\mathop{\boldsymbol{\rm{KNN}}}\limits_{k=k_{ms}}(\mathbf{P})),\mathcal{D}\in\{Lin,Pla\},$
where $\mathcal{D}_{k=k_{ms}}$ denotes the multi-scale structural domains, with $k_{ms}$ represents the set of scale parameters, and $\mathcal{S}(\cdot)$ refers to the process of calculating structural domains as previously described. In our study, we establish the default set of scale parameters as $\{10,20\}$, signifying that the linearity and planarity domains are computed using the 10-nearest-neighbor and 20-nearest-neighbor configurations, respectively.
3.2.3. Statistical Parameters Estimation
To quantify the quality representation from the structural domains, we employ some basic statistical parameters estimation process:
(9) $F_{S}=\{avg(\mathcal{D}_{k=k_{ms}}),$ $std(\mathcal{D}_{k=k_{ms}}),ent(\mathcal{D}_{k=k_{ms}})\},$
$\mathcal{D}\in\{Lin,Pla\},$
where $avg(\cdot)$, $std(\cdot)$, and $ent(\cdot)$ represent the average function, standard deviation function, and entropy function respectively, and $F_{S}$ indicates the set of the final extracted structural features.
3.3. Quality Regression
To clearly demonstrate the efficacy of the proposed features, we merge the LMM evaluation results with the structural features, and then incorporate them into the visual quality score using support vector regression (SVR):
(10) $Q=\mathbf{SVR}(F_{L}\oplus F_{S}),$
where $Q$ indicates the quality values, $\mathbf{SVR}(\cdot)$ represents the SVR regression process, and $\oplus$ denotes the concatenation process.
4. Experiment
4.1. Validation Databases
To assess the efficacy of the proposed method, we employ the SJTU-PCQA database (Yang et al., 2020a), the Waterloo point cloud assessment database (WPC)(Liu et al., 2022b), and the WPC2.0 database (Liu et al., 2021a) for validation. The SJTU-PCQA database contains 9 reference point clouds, subjected to seven distortion types (compression, color noise, geometric shift, down-sampling, and three mixed distortions) at six levels, yielding a total of 378 ($9\times 7\times 6$) distorted point clouds. The WPC database includes 20 reference point clouds, each modified by four distortions (down-sampling, Gaussian white noise, Geometry-based Point Cloud Compression (G-PCC), and Video-based Point Cloud Compression (V-PCC)), resulting in 740 ($20\times 37$) distorted point clouds. Meanwhile, the WPC2.0 database features 16 reference point clouds, each undergoing 25 different V-PCC degradation settings, leading to 400 ($16\times 25$) distorted point clouds.
Table 1. Performance on the SJTU-PCQA, WPC, and WPC2.0 databases. Best in red, second in blue.
Type Methods SJTU-PCQA WPC WPC2.0
[…]
FR MSE-p2po 0.7294 0.8123 0.5617 1.3613 0.4558 0.4852 0.3182 19.8943 0.4315 0.4626 0.3082 19.1605
HD-p2po 0.7157 0.7753 0.5447 1.4475 0.2786 0.3972 0.1943 20.8990 0.3587 0.4561 0.2641 18.8976
MSE-p2pl 0.6277 0.5940 0.4825 2.2815 0.3281 0.2695 0.2249 22.8226 0.4136 0.4104 0.2965 21.0400
HD-p2pl 0.6441 0.6874 0.4565 2.1255 0.2827 0.2753 0.1696 21.9893 0.4074 0.4402 0.3174 19.5154
PSNR-yuv 0.7950 0.8170 0.6196 1.3151 0.4493 0.5304 0.3198 19.3119 0.3732 0.3557 0.2277 20.1465
PCQM 0.8644 0.8853 0.7086 1.0862 0.7434 0.7499 0.5601 15.1639 0.6825 0.6923 0.4929 15.6314
GraphSIM 0.8783 0.8449 0.6947 1.0321 0.5831 0.6163 0.4194 17.1939 0.7405 0.7512 0.5533 14.9922
PointSSIM 0.6867 0.7136 0.4964 1.7001 0.4542 0.4667 0.3278 20.2733 0.4810 0.4705 0.2978 19.3917
[…]
NIQE 0.1379 0.2420 0.1009 2.2622 0.1136 0.2225 0.0953 23.1415 0.1865 0.2925 0.1335 22.5146
IL-NIQE 0.0837 0.1603 0.0594 2.3378 0.0913 0.1422 0.0853 24.0133 0.0911 0.1233 0.0714 23.9987
IT-PCQA 0.8651 0.8283 0.6430 1.1661 0.4870 0.4329 0.3006 19.8960 0.5661 0.5432 0.3477 18.7224
ResSCNN 0.8600 0.8100 - - - - - - 0.7500 0.7200 - -
PQA-net 0.8372 0.8586 0.6304 1.0719 0.7026 0.7122 0.4939 15.0812 0.6191 0.6426 0.4606 16.9756
3D-NSS 0.7144 0.7382 0.5174 1.7686 0.6479 0.6514 0.4417 16.5716 0.5077 0.5699 0.3638 17.7219
GMS-3DQA 0.9108 0.9177 0.7735 0.7872 0.8308 0.8338 0.6457 12.2292 0.8272 0.8218 0.6277 12.9904
MM-PCQA 0.9103 0.9226 0.7838 0.7716 0.8414 0.8556 0.6513 12.3506 0.8023 0.8024 0.6202 13.4289
LMM-PCQA(Ours) 0.9376 0.9404 0.8002 0.7175 0.8825 0.8739 0.7064 11.8171 0.8614 0.8634 0.6723 10.6924
4.2. Implementation Details
4.2.1. LMM Training
[…]
$\mathcal{H}_{t}$ $=\mathcal{E}_{t}(t),$
$\mathcal{H}$ $=\mathcal{H}_{v}\oplus\mathcal{H}_{t},$
$\mathcal{O}$ $=\mathcal{L}(\mathcal{H}),$
where $\mathcal{H}_{v}$ and $\mathcal{H}_{t}$ represent the abstracted tokens for the visual and text input respectively, $\mathcal{O}$ stands for the output. For all PCQA databases, the batch size is maintained at 64. The learning rate is fixed at $2\times 10^{-5}$, with the training process extending over 2 epochs for each variant. We utilize the common GPT (Radford et al., 2019) loss mechanism, specifically the cross-entropy between the predicted logits and actual labels. The evaluation of performance metrics is conducted using the final weights obtained post-training. Four NVIDIA A100 80G GPUs are employed for the training phase, while a single RTX3090 24G GPU is used to measure inference latency. In the inference stage, only the input texts preceding the [SCORE_TOKEN] are inputted into the LMM, leading to the final element of $\mathbf{O}$ representing the targeted probability map.
4.2.2. Validation Strategy
Following the methodologies in (Fan et al., 2022b; Zhang et al., 2023a; Chai et al., 2024), we utilize a k-fold cross-validation approach in our experiments to ensure a dependable performance evaluation of our proposed method. The SJTU-PCQA, WPC, and WPC2.0 databases consist of 9, 20, and 16 point cloud groups, respectively, leading us to adopt 9-fold, 5-fold, and 4-fold cross-validation for these databases to achieve an approximate 8:2 train-test split. The average of the performance metrics is considered the definitive result. It is crucial to note that the training and testing sets are mutually exclusive to prevent content overlap. For FR-PCQA methods, which do not require training, we evaluate them using the same test sets and report the average performance.
4.3. Competitors
17 state-of-the-art quality assessment methods are selected for comparison, which consists of 8 FR-PCQA and 9 NR-PCQA methods:
  • The FR-PCQA methods include MSE-p2point (MSE-p2po) (Mekuria et al., 2016b), Hausdorff-p2point (HD-p2po) (Mekuria et al., 2016b), MSE-p2plane (MSE-p2pl) (Tian et al., 2017), Hausdorff-p2plane (HD-p2pl) (Tian et al., 2017), PSNR-yuv (Torlig et al., 2018), PCQM (Meynet et al., 2020), GraphSIM (Yang et al., 2020b), and PointSSIM (Alexiou and Ebrahimi, 2020).
  • The NR-PCQA methods include BRISQUE (Mittal et al., 2012a), NIQE (Mittal et al., 2012b), IL-NIQE (Zhang et al., 2015b), IT-PCQA (Yang et al., 2022), ResSCNN (Liu et al., 2022c), PQA-net (Liu et al., 2021b), 3D-NSS (Zhang et al., 2022b), GMS-3DQA (Zhang et al., 2023b) and MM-PCQA (Zhang et al., 2023a).
Note that BRISQUE, NIQE, IL-NIQE are image-based quality assessment metrics and are validated on the same projections.
4.4. Evaluation Criteria
Four mainstream evaluation criteria in the quality assessment field are utilized to compare the correlation between the predicted scores and MOSs, which include Spearman Rank Correlation Coefficient (SRCC), Kendall’s Rank Correlation Coefficient (KRCC), Pearson Linear Correlation Coefficient (PLCC), Root Mean Squared Error (RMSE). An excellent quality assessment model should obtain values of SRCC, KRCC, PLCC close to 1 and RMSE to 0.
4.5. Performance Discussion
The performance comparison between the proposed LMM-PCQA and other PCQA competitors on the SJTU-PCQA, WPC, and WPC2.0 databases is illustrated in Table 1, from which we can draw several conclusions: 1) The LMM-PCQA demonstrates superior performance across all three PCQA databases, even outperforming FR-PCQA methods. For instance, the proposed LMM-PCQA surpasses the second-best NR-PCQA method by about 0.027 (against GMS-3DQA), 0.041 (against MM-PCQA), and 0.034 (against GMS-3DQA) on the SJTU-PCQA, WPC, and WPC2.0 databases from the SRCC values. This highlights LMM’s capability to effectively assimilate PCQA knowledge and apply it actively. The consistency in LMM-PCQA’s performance across various databases underscores its potential to set new baselines in the PCQA field. 2) All PCQA competitors generally perform better on the SJTU-PCQA database but face significant performance declines on the WPC and WPC2.0 databases. In contrast, LMM-PCQA exhibits the smallest performance drops, with decreases of approximately 0.06 and 0.08 in SRCC values when transitioning from the SJTU-PCQA to the WPC and WPC2.0 databases, respectively. This performance stability underscores LMM-PCQA’s robustness and superior ability to handle diverse content effectively, showcasing its adaptability and consistency in quality assessment across different point cloud databases.
Table 2. Contributions of LMM evaluation results and structural features, where ‘w/o LMM’ indicates excluding the LMM evaluation results, ‘w/o Structural’ indicates excluding the structural features.
Modal SJTU-PCQA WPC WPC2.0
[…]
w/o LMM 0.6650 0.7274 0.3598 0.3523 0.3847 0.3951
w/o Structural 0.9081 0.9158 0.8488 0.8271 0.8258 0.8381
LMM + Structural 0.9376 0.9404 0.8825 0.8739 0.8614 0.8634
Figure 4. SRCC/PLCC performance tendency according to the number of used projections on the SJTU-PCQA, WPC, and WPC2.0 databases.
4.6. Ablation Study
4.6.1. Contributions of LMM evaluation results and structural features
To fully investigate the contributions and validate the rationality behind the proposed dual streams of features, we decide to undertake an ablation study in this section. The results, as detailed in Table 2, clearly demonstrate that the integration of both feature streams leads to superior performance compared to employing a single feature stream. Upon a detailed examination, it is apparent that LMM evaluation results markedly surpass the structural features in terms of performance. This disparity primarily stems from the fact that some distorted point clouds within the SJTU-PCQA, WPC, and WPC2.0 databases suffer only from color distortions. Structural features, derived from geometric information, inherently lack the capability to recognize these color distortions, resulting in their relatively lower performance in total. However, incorporating structural features enhances the comprehensive understanding of point cloud quality, contributing to the refinement and elevation of the assessment precision in LMM evaluation results. This combined method highlights how important it is to look at features from different angles to fully understand and capture the subtle details of point cloud quality.
4.6.2. Influence of the number of projections
The proposed LMM-PCQA utilizes 6 projections as default. In this section, we further change the number of used projections to test the corresponding performance influence. Specifically, we randomly select 1-6 projections from the cube-like 6 projections setting as the input projections of LMM-PCQA. The performance tendency is illustrated in Fig. 4, from which we can draw several interesting conclusions: 1) As the number of projections increases, the performance of LMM-PCQA also improves correspondingly, indicating that increasing the number of projections can encompass more effective quality information, thereby enhancing the final performance. 2) Specifically, when the number of projections increases from 2 to 5, the improvement is significantly more pronounced. This suggests that at this stage, the quality information is not yet redundant, and the benefit of increasing the number of projections is relatively large. However, when the number of projections increases from 5 to 6, the performance improvement is relatively low, indicating that the quality information has become somewhat saturated, and further increases in the number of projections yield diminishing returns.
4.6.3. Effect of the multi-scale structural features
To quantify the contributions of the multi-scale mechanism, we validate the performance of structural features with different scale parameters under two settings: with LMM evaluation results and without LMM evaluation results. The experimental performance is listed in Table 4.6.3. From the table, we can find that the multi-scale structural features with k=10,20 perform better than the single-scale features whether the LMM evaluation results are involved or not, which confirms the effectiveness of the proposed multi-scale mechanism. This can be attributed to that humans tend to perceive the visual quality of point clouds from a multi-scale perspective.
Table 3. Performance of the multi-scale structural features, where k is the scale parameter of the KNN algorithm.
Model SJTU-PCQA WPC WPC2.0
[…]
w/o LMM
[…]
k=20 0.5920 0.6311 0.1795 0.2720 0.3224 0.3129
k=10,20 0.6650 0.7274 0.3598 0.3523 0.3847 0.3951
[…]
k=20 0.9199 0.9179 0.8466 0.8488 0.8578 0.8562
k=10,20 0.9376 0.9404 0.8825 0.8739 0.8614 0.8634
[…]
PQA-net 0.5411 0.6102 0.6006 0.6377
3D-NSS 0.1817 0.2344 0.4933 0.5613
GMS-3DQA 0.7421 0.7611 0.7822 0.7714
MM-PCQA 0.7991 0.7902 0.7917 0.7935
LMM-PCQA(Ours) 0.8246 0.7999 0.8385 0.8387
Table 5. Distortion-specific performance results on the SJTU-PCQA database, where OT represents octree-based compression, CN represents color noise, DS represents down-sampling, DS+CN represents down-sampling and color noise, DS+GGN represents down-sampling and geometry Gaussian noise, GGN represents geometry Gaussian noise, and CN+GGN represents color noise and geometry Gaussian noise respectively.
Distortion OT CN DS DS+CN DS+GGN GGN CN+GGN
[…]
MSE-p2po 0.71 0.76 nan nan 0.93 0.95 0.96 0.87 0.96 0.89 0.98 0.89 0.99 0.90
HD-p2po 0.64 0.69 nan nan 0.82 0.88 0.80 0.75 0.94 0.91 0.98 0.91 0.99 0.91
MSE-p2pl 0.55 0.62 nan nan 0.87 0.92 0.85 0.81 0.96 0.75 0.97 0.85 0.98 0.86
HD-p2pl 0.54 0.58 nan nan 0.82 0.87 0.81 0.79 0.94 0.77 0.95 0.88 0.97 0.85
PSNR-yuv 0.59 0.54 0.86 0.87 0.91 0.91 0.96 0.91 0.97 0.94 0.98 0.95 0.99 0.96
PCQM 0.80 0.84 0.86 0.85 0.93 0.96 0.97 0.94 0.96 0.90 0.98 0.93 0.99 0.93
GraphSIM 0.82 0.81 0.82 0.90 0.96 0.97 0.91 0.95 0.95 0.95 0.96 0.97 0.97 0.98
PointSSIM 0.80 0.88 0.87 0.87 0.93 0.93 0.97 0.93 0.96 0.97 0.98 0.97 0.99 0.96
PQA-net 0.81 0.82 0.84 0.83 0.91 0.92 0.93 0.91 0.89 0.89 0.95 0.96 0.97 0.96
3D-NSS 0.60 0.67 0.85 0.79 0.80 0.84 0.94 0.93 0.90 0.90 0.96 0.93 0.98 0.94
GMS-3DQA 0.83 0.84 0.91 0.92 0.95 0.95 0.95 0.93 0.96 0.97 0.97 0.93 0.98 0.95
MM-PCQA 0.84 0.83 0.92 0.93 0.94 0.96 0.94 0.95 0.94 0.93 0.95 0.94 0.98 0.96
LMM-PCQA 0.89 0.90 0.97 0.97 0.97 0.98 0.95 0.94 0.98 0.98 0.96 0.93 0.98 0.97
Table 6. Distortion-specific performance results on the WPC database, where DS represents down-sampling, GN represents geometry and color Gaussian noise, G-PCC represents geometry-based compression, and V-PCC represents video-based compression.
Distortion DS GN G-PCC V-PCC
[…]
MSE-p2po 0.46 0.46 0.67 0.63 0.84 0.72 0.41 0.44
HD-p2po 0.39 0.35 0.69 0.70 0.80 0.70 0.36 0.36
MSE-p2pl 0.41 0.40 0.61 0.47 0.61 0.48 0.42 0.40
HD-p2pl 0.40 0.41 0.62 0.51 0.57 0.59 0.41 0.41
PSNR-yuv 0.23 0.28 0.79 0.88 0.47 0.43 0.44 0.48
PCQM 0.66 0.64 0.89 0.89 0.86 0.77 0.83 0.78
GraphSIM 0.56 0.57 0.79 0.81 0.75 0.74 0.71 0.70
PointSSIM 0.35 0.34 0.83 0.87 0.91 0.95 0.51 0.41
PQA-net 0.61 0.63 0.77 0.78 0.87 0.88 0.76 0.77
3D-NSS 0.55 0.51 0.81 0.83 0.86 0.87 0.49 0.47
GMS-3DQA 0.72 0.73 0.87 0.88 0.89 0.89 0.86 0.88
MM-PCQA 0.75 0.74 0.88 0.87 0.88 0.89 0.89 0.86
LMM-PCQA 0.76 0.79 0.91 0.91 0.96 0.96 0.90 0.89
4.7. Cross-database Validation
[…]
4.8. Distortion-specific Evaluation
To verify the effectiveness of the proposed LMM-PCQA on different kinds of distortions, we carry out the distortion-specific evaluation experiment in this section. The performance comparison is shown in Table 5 (The ‘nan’ values for MSE-p2po, HD-p2po, MSE-p2pl, and HD-p2pl arise because these methods only analyze the geometric differences between the reference and distorted point clouds, thereby failing to account for color distortions.) and Table 6, from which we can obtain several findings: 1) The proposed LMM-PCQA achieves the best performance on 4 of 7 distortion types and all distortion types on the SJTU-PCQA database and the WPC database respectively, which suggests that LMM-PCQA is effective at dealing with various kinds of distortions. 2) Although LMM-PCQA does not achieve the top performance on DS+CN, GGN, and CN+GGN distortions within the SJTU-PCQA database, the performance gap to the best is minimal, with a difference of no more than 0.02 in terms of SRCC values. This suggests that LMM-PCQA remains in the top tier. 3) Upon examining Table 6 more closely, it is evident that all PCQA methods exhibit substantial declines in performance with the ‘DS’ distortion. The underlying reason for this trend is the simplicity of the point cloud reference models used in the WPC database, leading to a reduced sensitivity to downsampling distortion.
5. Conclusion
In conclusion, this paper pioneers the integration of LMMs with PCQA, unveiling the untapped potential of LMMs in this domain. Our research successfully demonstrates the feasibility of adapting LMMs for PCQA through text supervision, enhancing their capability to evaluate 3D visual quality from 2D projections. We also propose to capture multi-scale structural features to offer a more holistic view of the point cloud quality. By combining LMM evaluation results and structural features, our approach significantly improves the accuracy of PCQA. We carry out thorough experiments to demonstrate the effectiveness of the proposed LMM-PCQA, as well as its robustness and ability to generalize across various distortion types and diverse point cloud content. The encouraging results not only validate our methodology but also lay the groundwork for future research. We hope our work will serve as a stepping stone for further exploration into the synergy between LMMs and PCQA, driving forward innovations in 3D visual quality analysis.
Acknowledgements.
This work was supported in part by the National Natural Science Foundation of China (623B2073, 62101326, 62225112, 62301316), the Foundation of Key Laboratory of Media Audio & Video (Communication University of China), Ministry of Education, China (JGKFKT2302), the China Postdoctoral Science Foundation under Grants 2023TQ0212 and 2023M742298, and the Postdoctoral Fellowship Program of CPSF under Grant GZC20231618.